Evening shift August 6, 1999

B. Cox (cox@fnal.gov)
Sat, 7 Aug 1999 07:45:46 -0500 (CDT)

Shift Summary for Evening Shift, 7/8/99

On Shift: Brad, Theo

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 6.0
Hours of E832 data (special runs not included): 7.0
Hours of special runs (laser scans, muon runs...): 0.

Accounting of lost time:

Crate 53 work 15:30 - 16:05
RF Problems 19:32 - 20:25
Tape change 21:30 - 21:40
========================================================================
Runs ended this shift:

run 14131
run 14132

=======================================================================

Tapes taken to FCC this shift

QK*222-224, QKB078
QK*225-227 QKB079

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

Reasonably steady running. A few kfrends No real problems

========================================================================
Description of Shift:

15:40 - 16:05 latch in slot 13 crate 53 removed since it
was supposed to be the source of all the
ktevana errors and Kfrend requests

16:05 Started running again on run 14131

16:08 Beam goes away
16:25 Beam returns
19:32 Beam off for RF problems
20:25 Beam returns
21:30 Tapes QK*225-227 finish - Mount new tapes QK*228-230
21:40 Start new run 14133

21:40-34:00 Quite a few empty spills - roughly 60-70% efficient

========================================================================
Follow up Notes on Previous Shift

Keep taking data. No magnet flip, no muon run

If opportunity arises (beam off for a period), take a laser scan.